# DeltaEdit: Enhancing Sequential Editing in Large Language Models by Controlling Superimposed Noise

### v1

@@ 4.2 The Implementation of DeltaEdit
$noise_{e}=\|\sum_{i\leq e}\Delta_{i}k_{e}\|_{2}^{2}-\|\Delta_{e}k_{e}\|_{2}^{2}=\|\sum_{i\leq e-1}\Delta_{i}k_{e}\|_{2}^{2}+2\sum_{i\leq e-1}k_{e}^{\top}\beta_{e}\alpha_{e}^{\top}\alpha_{i}\beta_{i}^{\top}k_{e}$ (16)
This formula shows that $noise_{e}$ is determined by the cumulative interference from historical edits $(\|\sum_{i\leq e-1}\Delta_{i}k_{e}\|_{2}^{2})$ and the interaction between the current and historical edits $(\sum_{i\leq e}k_{e}^{\top}\beta_{e}\alpha_{e}^{\top}\alpha_{i}\beta_{i}^{\top}k_{e})$. It can be observed that the latter term can be reduced by constraining the training of $\alpha$.
@@
$\Delta_{\text{history}}=\sum_{i\leq e-1}\Delta_{i},$ (17)
@@
To achieve orthogonality between the $\alpha_{e}$ and all vectors in the set $A=\{\alpha_{i},i<e\}$ without storing these vectors, we compute a null space[35] using singular value decomposition (SVD). Restricting the training of $\alpha_{e}$ to this null space ensures that $\alpha_{e}$ remains almost orthogonal to all vectors in $A$, effectively avoiding the storage overhead. To compute this null space, the column space matrix $D$ is first constructed:
@@
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of DeltaEdit. First, we assess its performance in sequential editing tasks. Subsequently, we evaluate its effectiveness in mitigating $noise_{E}$ and examine the model’s generalization capabilities after the editing process.
@@ 5.1 Experiment Setup
In experiments, we adopt the computation method from AlphaEdit to calculate $\beta$. The experiments are conducted on two language models: GPT2-XL[33] and Llama3-8B[34]. To comprehensively evaluate the effectiveness of DeltaEdit, we compare its performance with several baseline methods, including Fine-Tuning[36], ROME[10], MEMIT[20], PRUNE[22], RECT[23], and AlphaEdit[21].
The evaluation is conducted on two widely recognized benchmark datasets: ZsRE[37] and CounterFact[10]. Following prior work[20], for ZsRE, we employ Efficacytop, Generalizationtop, and Specificitytop as evaluation metrics. For CounterFact, we employ Efficacylarger, Generalizationlarger, and Specificitylarger. Additionally, for more comprehensive evaluation, Efficacytop, Generalizationtop, and Specificitytop are also employed. Metrics with the subscript “top” focus on the token with the highest output probability, while metrics with the subscript “larger” focus on the relative probability of the target token. Detailed definitions are provided in the appendix A. The experiments assess the performance of various methods after 3,000 sequential edits, highlighting the models’ abilities in editing efficiency, semantic generalization, and preserving unedited content.
@@ 5.2 Results of Experiment
DeltaEdit Demonstrates Superior Editing Performance  DeltaEdit demonstrates superior performance across most evaluation metrics compared to all baseline methods, underscoring its effectiveness in handling complex sequential editing tasks. Its performance is particularly strong on the CounterFact dataset. On the Llama3-8B model, DeltaEdit achieves 43.52%, 26.29%, and 15.94% higher scores than AlphaEdit in Efficacytop, Generalizationtop, and Specificitytop, respectively. On the GPT2-XL model, the improvement of DeltaEdit over AlphaEdit is smaller. This is because AlphaEdit already performs well on GPT2-XL, leaving less room for improvement. However, DeltaEdit still achieves better overall results. On the ZsRE dataset, DeltaEdit does not show a significant advantage. The lower similarity between object in ZsRE reduces the impact of superimposed noise during sequential editing. This suggests that DeltaEdit’s noise control mechanisms are particularly effective in scenarios with high object similarity, as seen in the CounterFact dataset.
DeltaEdit Enhances Stability in Sequential Editing  Sequentially performing 3,000 edits is an extremely challenging task, as frequent parameter updates can introduce more superimposed noise and degrade model performance over time. This issue is particularly evident in baseline methods, which perform poorly in this setting, especially on the Llama3-8B model. Despite its larger size, Llama3-8B has smaller parameter value ranges, making it more sensitive to noise accumulation. In contrast, DeltaEdit demonstrates remarkable stability and robustness during sequential editing. It maintains consistently strong performance, even after 3,000 consecutive edits, and significantly outperforms all baseline methods. These results highlight the effectiveness of DeltaEdit’s noise control strategy, which effectively suppresses noise growth to maintain high accuracy and stability in long-term editing tasks.
@@ 5.3 The Impact on Superimposed Noise
Figure 4(a) illustrates the changes in Efficacytop and $\textbf{noise}_{\textbf{E}}$ for GPT2-XL and Llama3-8B on the CounterFact dataset as the number of edits increases. As shown in these figures, DeltaEdit effectively reduces $noise_{E}$ while maintaining the stability of Efficacytop. This observation further reinforces the empirical support for the hypothesis that superimposed noise has a significant impact on the effectiveness of editing methods. Furthermore, it confirms that reducing superimposed noise can substantially enhance editing performance.
@@
5.4 General Capability Tests
To evaluate the impact on the general capability of the edited models, we conduct tests on six tasks from the General Language Understanding Evaluation (GLUE) benchmark[38]: CoLA[39], MMLU[40], MRPC[41], NLI[42], RTE[43], and SST[44]. The experimental results are shown in Figure 3, which depicts the F1 scores of the models and their corresponding F1 differences (i.e., the difference between the F1 scores of the edited and original models) as the number of edits increases on CounterFact.
(a) The changes in Efficacytop and $noise_{E}$ with the number of edits for GPT2-XL (top) and LLaMA3-8B (bottom).
(b) The distribution of hidden representations of pre-edited and post-edited Llama3-8B after dimensionality reduction. The top and right curve graphs display the marginal distributions for two reduced dimensions. The dashed lines represent the 0.95 confidence intervals.
Figure 4: Combined visualizations. (a) The changes in Efficacytop and $noise_{E}$ with the number of edits. (b) The distribution of hidden representations for Llama3-8B.
@@ 5.5 Hidden Representations Analysis
Knowledge editing can disrupt the distribution of the model’s hidden representations. To investigate the differences in representation distributions before and after editing, we randomly select 1,000 factual prompts. Hidden representations are then extracted from three versions of the Llama3-8B: the pre-edited model, the model edited using AlphaEdit, and the model edited using DeltaEdit. Both models were edited 3000 times on CounterFact. Finally, t-SNE is applied to the representations for dimensionality reduction to facilitate visualization. The results presented in Figure 4(b) show the reduced-dimensional distributions of the representations and their marginal distribution curves.
As shown in Figure 4(b), DeltaEdit is able to maintain the consistency of hidden representations before and after editing, whereas AlphaEdit causes significant shifts in hidden representations. We suggest that this minimal shift in representations is the critical factor enabling DeltaEdit to effectively maintain the model’s general capabilities. Notably, while DeltaEdit is originally designed to optimize representation shifts specifically for the edited tokens, it also demonstrates the ability to minimize post-editing shifts across all hidden representations. These results demonstrate that DeltaEdit successfully reduces the overall superimposed noise within the model.
@@ 6 Conclusion
In this work, we identify superimposed noise as a key challenge in sequential knowledge editing for large language models. To address this issue, we analyze two factors that influence the magnitude of superimposed noise and propose a novel method, DeltaEdit. This method employs a dynamic orthogonal constraint strategy to optimize update parameters. Our experiments on GPT2-XL and Llama3-8B demonstrate that DeltaEdit consistently outperforms existing methods in terms of edit success rates and the retention of general model capabilities. Additionally, our analysis of superimposed noise and hidden representations confirms that DeltaEdit preserves the integrity of the model’s internal states, ensuring stable and precise even under extensive sequential editing.
Overall, DeltaEdit offers a robust solution for maintaining the accuracy and reliability of large language models in dynamic environments, paving the way for more effective and scalable sequential editing techniques.
@@ A.1 Datasets
ZsRE[37] is a question answering dataset that uses questions generated through back-translation as equivalent neighboring questions. According to previous studies, Natural Questions are used as out-of-scope data to evaluate locality. Each sample in ZsRE includes a subject string and an answer as the editing target to evaluate editing success, along with paraphrased questions for generalization evaluation and locality questions for specificity evaluation.
Counterfact[10] is a more challenging dataset designed to compare counterfactual and factual statements. Each record in the dataset is derived from corresponding entries in PARAREL[45], where all subjects, relations, and objects are entities from WikiData. Out-of-scope data is constructed by replacing subject entities with approximate entities that share the same predicate. The Counterfact dataset uses metrics similar to those of ZsRE to evaluate efficacy, generalization, and specificity.
@@ A.3 Implementation Details
We implement DeltaEdit on GPT2-XL and Llama3-8B, following the configuration described in AlphaEdit[21]. The details are as follows:
@@
The curves in Figure 1, Figure 2 and Figure 4(a) are derived from the final edited layers of the models, specifically the 17th layer of GPT2-XL and the 8th layer of Llama3-8B, which are the layers most significantly impacted by the edits.
To enhance the experimental challenge, the batch size for each edit is set to 1 in all experiments in this paper.
@@ Appendix B Algorithm
12:        if Number of eigenvalues in $N$ exceeds $3/4$ of $\dim(\alpha_{e})$ then
13:           Remove smallest eigenvalues from $N$ until $|N|\leftarrow 3/4\times\dim(\alpha_{e})$
@@
15:        Select eigenvectors corresponding to remaining eigenvalues in $N$ to form $\hat{U}$
@@
[…]
26:        Update $\alpha_{e}$ by optimizing loss function (Equation 7)
@@ Appendix C Supplementary Experiments
C.1 Supplementary experimental results of Llama3-8B
In this section, we present the experimental results for editing Llama3-8B that are omitted in Sections 3.3 and 4.1 due to excessively large numerical values.
Table 2 reports the $noise_{E}$ values after applying MEMIT to edit Llama3-8B across different numbers of edits on CounterFact dataset. As shown, $noise_{E}$ increases rapidly to extremely large values, which explains MEMIT’s poor performance in Table 1. It is worth noting that while MEMIT achieves the highest Specificitylarger after editing Llama3-8B, this is primarily because most of its edits fail, resulting in little to no change in the output probability of the target tokens.
@@ C.1 Supplementary experimental results of Llama3-8B
C.2 The Impact of DeltaEdit on noiseE Across All Edited Layers
Figures 5 and 6 show the $noise_{E}$ in the multi-layer editing setup[20; 21] after applying AlphaEdit and DeltaEdit to edit GPT2-XL and Llama3-8B on CounterFact across different numbers of edits. Experimental results show that DeltaEdit can effectively reduce the $noise_{E}$ across all edited layers.
C.3 Analysis experiments of GPT2-XL
In this section, we present the results of the general capability tests and the hidden representations analysis of GPT2-XL.
Figure 7 demonstrates the test results of GPT2-XL on six tasks, edited using AlphaEdit and DeltaEdit on CounterFact. These six tasks are introduced in Section 5.4. It is evident that GPT2-XL has minimal ability to perform these tasks. Without any edits, GPT2-XL’s performance on five tasks (excluding SST) is close to random guessing, resulting in minimal performance changes after editing. On SST, although the performance after 3000 edits with DeltaEdit is lower than that with AlphaEdit, AlphaEdit appears to cause greater performance fluctuations compared to DeltaEdit.
Figure 8 presents the t-SNE dimensionality reduction results of GPT2-XL’s hidden representations after 3000 edits using AlphaEdit and DeltaEdit on CounterFact. It is evident that neither AlphaEdit nor DeltaEdit significantly alters the distribution of the hidden representations.
C.4 Impact of the Hyperparameter $\eta$
In this section, we demonstrate the impact of the hyperparameter $\eta$. Figure 9 presents the test results after 3000 edits on CounterFact using varying hyperparameter values of $\eta$. On GPT2-XL, $\text{Efficacy}_{\text{top}}$ shows minimal variation across different $\eta$ values. However, $\text{Generalization}_{\text{top}}$ rises significantly as $\eta$ increases from 2 to 2.5, while $\text{Specificity}_{\text{top}}$ drops sharply. Beyond $\eta=2.5$, these metrics stabilize. On Llama3-8B, as $\eta$ increases, both $\text{Efficacy}_{\text{top}}$ and $\text{Specificity}_{\text{top}}$ decrease, while $\text{Generalization}_{\text{top}}$ is notably higher at $\eta=2$ compared to other values. These results suggest that increasing $\eta$ beyond a certain threshold significantly affects editing performance: $\text{Generalization}_{\text{top}}$ improves, while $\text{Specificity}_{\text{top}}$ declines. Interestingly, $\text{Efficacy}_{\text{top}}$ exhibits contrasting trends across models: it increases on GPT2-XL but decreases on Llama3-8B. We hypothesize that this difference arises because small changes in $\eta$ lead to significant variations in the number of orthogonal strategy executions, which impacts performance. Additionally, the hidden representations of GPT2-XL have a lower dimensionality compared to those of Llama3-8B. This suggests that for GPT2-XL, excessive orthogonal intensity caused by small $\eta$ values may degrade editing performance. Above all, it is necessary to find an appropriate $\eta$ to achieve a balance between editing effectiveness, generalization capability, and the retention of other knowledge.

### v2

@@ 4 DeltaEdit
$noise_{e}$ $=\|\sum_{i\leq e}\Delta_{i}k_{e}\|_{2}^{2}-\|\Delta_{e}k_{e}\|_{2}^{2}$ (16)
$=\|\sum_{i<e}\Delta_{i}k_{e}\|_{2}^{2}+2\sum_{i<e}k_{e}^{\top}\beta_{e}\alpha_{e}^{\top}\alpha_{i}\beta_{i}^{\top}k_{e}$
This formula shows that $noise_{e}$ is determined by the cumulative interference from historical edits $(\|\sum_{i<e}\Delta_{i}k_{e}\|_{2}^{2})$ and the interaction between the current and historical edits $(\sum_{i<e}k_{e}^{\top}\beta_{e}\alpha_{e}^{\top}\alpha_{i}\beta_{i}^{\top}k_{e})$. It can be observed that the latter term can be reduced by constraining the training of $\alpha$.
@@
$\Delta_{\text{history}}=\sum_{i<e}\Delta_{i},$ (17)
@@
To achieve orthogonality between the $\alpha_{e}$ and all vectors in the set $A=\{\alpha_{i},i<e\}$ without storing these vectors, we compute a null space (wang2021training) using singular value decomposition (SVD). Restricting the training of $\alpha_{e}$ to this null space ensures that $\alpha_{e}$ remains almost orthogonal to all vectors in $A$, effectively avoiding the storage overhead. To compute this null space, the column space matrix $D$ is first constructed:
@@
[…]
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of DeltaEdit. First, we assess its performance in sequential editing tasks. Subsequently, we evaluate its effectiveness in mitigating $noise_{E}$ and its ability to maintain the model’s hidden representations.
@@ 5.1 Experiment Setup
In experiments, we adopt the computation method from AlphaEdit to calculate $\beta$. The experiments are conducted on two language models: GPT2-XL (Radford2019LanguageMA) and Llama3-8B (dubey2024llama). To comprehensively evaluate the effectiveness of DeltaEdit, we compare its performance with several baseline methods, including Fine-Tuning (zhu2020modifying), ROME (meng2023locatingeditingfactualassociations), MEMIT (meng2023masseditingmemorytransformer), PRUNE (ma2024perturbationrestrainedsequentialmodelediting), RECT (gu2024modeleditingharmsgeneral), and AlphaEdit (fang2024alphaedit).
The evaluation is conducted on two widely recognized benchmark datasets: ZsRE (levy-etal-2017-zero) and CounterFact (meng2023locatingeditingfactualassociations). Following prior work (meng2023masseditingmemorytransformer), for ZsRE, we employ Efficacytop, Generalizationtop, and Specificitytop as evaluation metrics. For CounterFact, we employ Efficacylarger, Generalizationlarger, and Specificitylarger. Additionally, for more comprehensive evaluation, Efficacytop, Generalizationtop, and Specificitytop are also employed. Metrics with the subscript “top” focus on the token with the highest output probability, while metrics with the subscript “larger” focus on the relative probability of the target token. Detailed definitions are provided in the Appendix A.2. The experiments assess the performance of various methods after 3,000 sequential edits, highlighting the models’ abilities in editing efficiency, semantic generalization, and preserving unedited content.
@@ 5.2 Results of Experiment
DeltaEdit Demonstrates Superior Editing Performance. DeltaEdit demonstrates superior performance across most evaluation metrics compared to all baseline methods, underscoring its effectiveness in handling complex sequential editing tasks. Its performance is particularly strong on the CounterFact dataset. On the Llama3-8B model, DeltaEdit achieves 16.8%, 2.6%, and 16.61% higher scores than AlphaEdit in Efficacytop, Generalizationtop, and Specificitytop, respectively. On the GPT2-XL model, the improvement of DeltaEdit over AlphaEdit is smaller. This is because AlphaEdit already performs well on GPT2-XL, leaving less room for improvement. However, DeltaEdit still achieves better overall results. On the ZsRE dataset, DeltaEdit does not show a significant advantage. The lower similarity between object in ZsRE reduces the impact of superimposed noise during sequential editing. This suggests that DeltaEdit’s noise control mechanisms are particularly effective in scenarios with high object similarity, as seen in the CounterFact dataset.
Figure 4: The changes in Efficacytop and $noise_{E}$ with the number of edits. The top figure shows results for GPT2-XL, the bottom figure shows results for LLaMA3-8B.
DeltaEdit Enhances Stability in Sequential Editing. Sequentially performing 3,000 edits is an extremely challenging task, as frequent parameter updates can introduce more superimposed noise and degrade model performance over time. This issue is particularly evident in baseline methods, which perform poorly in this setting, especially on the Llama3-8B model. Despite its larger size, Llama3-8B has smaller parameter value ranges, making it more sensitive to noise accumulation. In contrast, DeltaEdit demonstrates remarkable stability and robustness during sequential editing. It maintains consistently strong performance, even after 3,000 consecutive edits, and significantly outperforms all baseline methods. These results highlight the effectiveness of DeltaEdit’s noise control strategy, which effectively suppresses noise growth to maintain high accuracy and stability in long-term editing tasks.
@@ 5.3 The Impact on Superimposed Noise
Figure 4 illustrates the changes in Efficacytop and $noise_{E}$ for GPT2-XL and Llama3-8B on the CounterFact dataset as the number of edits increases. As shown in these figures, DeltaEdit effectively reduces $\textbf{{noise}}_{\textbf{{E}}}$ while maintaining the stability of Efficacytop. This observation further reinforces the empirical support for the hypothesis that superimposed noise has a significant impact on the effectiveness of editing methods. Furthermore, it confirms that reducing superimposed noise can substantially enhance editing performance.
@@
Figure 5: The distribution of hidden representations of pre-edited and post-edited Llama3-8B after dimensionality reduction. The top and right curve graphs display the marginal distributions for two reduced dimensions. The dashed lines represent the 0.95 confidence intervals.
@@ 5.4 Hidden Representations Analysis
Knowledge editing can alter a model’s hidden representations. This study investigates these changes by comparing hidden representations from three Llama3-8B versions: pre-edited, edited with AlphaEdit, and edited with DeltaEdit (both edited 3,000 times on CounterFact). Representations were extracted using 1,000 factual prompts and visualized via t-SNE. Results in Figure 5 indicate that DeltaEdit largely preserves the original distribution of hidden representations, unlike AlphaEdit, which causes significant shifts. This minimal representational shift by DeltaEdit is crucial for maintaining the model’s general capabilities.
@@ 6 Conclusion
In this work, we identify superimposed noise accumulation as a critical limitation in sequential editing of large language models. Comprehensive analysis demonstrates that this phenomenon leads to degraded editing performance and model collapse. Additionally, we trace this to the interference of correct knowledge by erroneously activated irrelevant knowledge.
Based on our analysis, we propose DeltaEdit, which incorporates dynamic orthogonal constraints to mitigate superimposed noise by optimizing influence vectors that existing methods neglect. Extensive experiments show that DeltaEdit achieves superior sequential editing performance with 16.8% improvement over strong baselines while better preserving model capabilities and representation integrity, thus providing a robust solution for reliable sequential knowledge editing.
@@ A.1 Datasets
  • ZsRE (levy-etal-2017-zero) is a question answering dataset that uses questions generated through back-translation as equivalent neighboring questions. According to previous studies, Natural Questions are used as out-of-scope data to evaluate locality. Each sample in ZsRE includes a subject string and an answer as the editing target to evaluate editing success, along with paraphrased questions for generalization evaluation and locality questions for specificity evaluation.
  • Counterfact (meng2023locatingeditingfactualassociations) is a more challenging dataset designed to compare counterfactual and factual statements. Each record in the dataset is derived from corresponding entries in PARAREL (elazar2021measuring), where all subjects, relations, and objects are entities from WikiData. Out-of-scope data is constructed by replacing subject entities with approximate entities that share the same predicate. The Counterfact dataset uses metrics similar to those of ZsRE to evaluate efficacy, generalization, and specificity.
@@ A.3 Implementation Details
We implement DeltaEdit on GPT2-XL and Llama3-8B, following the configuration described in AlphaEdit (fang2024alphaedit). All experiments are conducted on Linux system and implemented using PyTorch. The details are as follows:
@@
The curves in Figure 2, Figure 3 and Figure 4 of the main paper are derived from the final edited layers of the models, specifically the 17th layer of GPT2-XL and the 8th layer of Llama3-8B, which are the layers most significantly impacted by the edits.
To enhance the experimental challenge, the batch size for each edit is set to 1 in all experiments in this paper. Our experimental results in Table 1 of the main paper are the average values obtained from three repeated experiments.
@@ Appendix B Algorithm
12:   if Number of eigenvalues in NN exceeds $3/4$ of $\dim(\alpha_{e})$ then
13:    Remove smallest eigenvalues from NN until $|N|\leftarrow 3/4\times\dim(\alpha_{e})$
@@
15:   Select eigenvectors corresponding to remaining eigenvalues in NN to form $\hat{U}$
@@
25:  for $\text{n}\in\text{num\_teps}$ do
26:   Update $\alpha_{e}$ by optimizing loss function
@@ Appendix C Supplementary Experiments
C.1 General Capability Tests
To evaluate the impact on the general capability of the edited models, we conduct tests on six tasks from the General Language Understanding Evaluation (GLUE) benchmark(wang2018glue): CoLA(warstadt2019neural), MMLU(hendrycks2020measuring), MRPC(dolan2005automatically), NLI(williams2017broad), RTE(bentivogli2009fifth), and SST(socher2013recursive). The experimental results on Llama3-8B are shown in Figure 6, which depicts the F1 scores of the models and their corresponding F1 differences (i.e., the difference between the F1 scores of the edited and original models) as the number of edits increases on CounterFact.
Figure 7 demonstrates the test results of GPT2-XL on six tasks, edited using AlphaEdit and DeltaEdit on CounterFact. It is evident that GPT2-XL has minimal ability to perform these tasks. Without any edits, GPT2-XL’s performance on five tasks (excluding SST) is close to random guessing, resulting in minimal performance changes after editing. On SST, although the performance after 3,000 edits with DeltaEdit is lower than that with AlphaEdit, AlphaEdit appears to cause greater performance fluctuations compared to DeltaEdit.
C.2 Supplementary experimental results of Llama3-8B
In this section, we present the experimental results for editing Llama3-8B that are omitted in Section 3.3 and Section 3.4 due to excessively large numerical values.
Table 2 reports the $noise_{E}$ values after applying MEMIT to edit Llama3-8B across different numbers of edits on CounterFact dataset. As shown, $noise_{E}$ increases rapidly to extremely large values, which explains MEMIT’s poor performance in Table 1 of the main paper. It is worth noting that while MEMIT achieves the highest Specificitylarger after editing Llama3-8B, this is primarily because most of its edits fail, resulting in little to no change in the output probability of the target tokens.
@@ C.2 Supplementary experimental results of Llama3-8B
C.3 The Impact of DeltaEdit on noiseE Across All Edited Layers
Figures 8 and 9 show the $noise_{E}$ in the multi-layer editing setup(meng2023masseditingmemorytransformer; fang2024alphaedit) after applying AlphaEdit and DeltaEdit to edit GPT2-XL and Llama3-8B on CounterFact across different numbers of edits. Experimental results show that DeltaEdit can effectively reduce the $noise_{E}$ across all edited layers.
C.4 Analysis experiments of GPT2-XL
In this section, we present the results of the hidden representations analysis of GPT2-XL.
Figure 10 presents the t-SNE dimensionality reduction results of GPT2-XL’s hidden representations after 3,000 edits using AlphaEdit and DeltaEdit on CounterFact. It is evident that neither AlphaEdit nor DeltaEdit significantly alters the distribution of the hidden representations.
C.5 Impact of the Hyperparameter $\eta$
In this section, we demonstrate the impact of the hyperparameter $\eta$. Figure 11 presents the test results after 3,000 edits on CounterFact using varying hyperparameter values of $\eta$. On GPT2-XL, $\text{Efficacy}_{\text{top}}$ shows minimal variation across different $\eta$ values. However, $\text{Generalization}_{\text{top}}$ rises significantly as $\eta$ increases from 2 to 2.5, while $\text{Specificity}_{\text{top}}$ drops sharply. Beyond $\eta=2.5$, these metrics stabilize. On Llama3-8B, as $\eta$ increases, both $\text{Efficacy}_{\text{top}}$ and $\text{Specificity}_{\text{top}}$ decrease, while $\text{Generalization}_{\text{top}}$ is notably higher at $\eta=2$ compared to other values. These results suggest that increasing $\eta$ beyond a certain threshold significantly affects editing performance: $\text{Generalization}_{\text{top}}$ improves, while $\text{Specificity}_{\text{top}}$ declines. Interestingly, $\text{Efficacy}_{\text{top}}$ exhibits contrasting trends across models: it increases on GPT2-XL but decreases on Llama3-8B. We hypothesize that this difference arises because small changes in $\eta$ lead to significant variations in the number of orthogonal strategy executions, which impacts performance. Additionally, the hidden representations of GPT2-XL have a lower dimensionality compared to those of Llama3-8B. This suggests that for GPT2-XL, excessive orthogonal intensity caused by small $\eta$ values may degrade editing performance. Above all, it is necessary to find an appropriate $\eta$ to achieve a balance between editing effectiveness, generalization capability, and the retention of other knowledge.